# FiDeLiS: Faithful Reasoning in Large Language Model for Knowledge Graph Question Answering

### v1

@@ 5 Conclusion
Appendix A Defintions
A.1 Defining a Valid Reasoning Trace/Path
Definition 1. A reasoning step is a pair $(r,e)$, where $r$ is the relation and $e$ is the corresponding entity.
Definition 2. A reasoning path $\mathcal{P}$ is a pair $(s,\mathcal{T})$, where $s$ is the starting entity for the reasoning path, and $\mathcal{T}$ is a sequence of reasoning steps $\mathcal{T}=\left\{t_{0},t_{1},\ldots,t_{n}\right\}$ and $t_{k}=(r_{k},e_{k})$. The $t_{k}\in\mathcal{R}$ denotes the $k$-th reasoning step in the path and $n$ denotes the length of the path.
[…]
A.2 Knowledge Graph-based Question Answering (KGQA)
@@
Knowledge Graph Question Answering (KGQA) is a typical reasoning task based on KGs. Given a natural language question $q$ and a $\mathrm{KG}$ $\mathcal{G}$, the task aims to design a function $f$ to predict answers $a\in\mathcal{A}_{q}$ based on knowledge from $\mathcal{G}$, i.e., $a=f(q,\mathcal{G})$. Following previous works [29, 28], we assume the entities $e_{q}\in\mathcal{T}_{q}$ mentioned in $q$ and answers $a\in\mathcal{A}_{q}$ are labeled and linked to the corresponding entities in $\mathcal{G}$, i.e., $\mathcal{T}_{q},\mathcal{A}_{q}\subseteq\mathcal{E}$.
Appendix B Prompt List
B.1 Plan-and-solve
In-Context Few-shot
Q: {Query}
A:
B.2 Deductive-verification
In-Context Few-shot
A:
B.3 Adequacy-verification
In-Context Few-shot
A:
B.4 Beam Search
In-Context Few-shot
A:
B.5 Reasoning
In-Context Few-shot
Question: {Query}
Reasoning path: {reasoning_path}
Only return the answer to the question.
A:
Appendix C Potential Impacts and Limitations
@@
The proposed retrieval-exploration interactive method, FiDeLiS, holds the potential to significantly enhance the performance of LLMs by addressing the issue of hallucinations and enabling deep, responsible reasoning. By integrating KGs with LLMs, the method offers more accurate knowledge retrieval, leveraging deductive reasoning capabilities to guide the reasoning process and avoid logical fallacies. The stepwise and generalizable approach, along with the use of deductive verification as a stop criterion, may reduce unnecessary computation and misleading reasoning chains. Additionally, the training-free nature of this method combined with its lower computational cost could be easily acted as a plugin-in for other methods/frameworks.
@@
However, there are limitations to consider. As the method is still in its early stages of development, there may be unforeseen challenges or drawbacks when applied to real-world scenarios. The reliance on external knowledge graphs means that the quality and coverage of these resources can impact the overall effectiveness of the method. There may also be computational constraints when working with very large or complex graphs. Additionally, while the method has shown promise in benchmarks, its performance in more diverse or specialized tasks is yet to be evaluated. Despite these challenges, the research presents a promising direction for improving the reasoning capabilities of LLMs.
Appendix D Efficiency of FiDeLiS
@@
The bottleneck of computation is the beam search process, which contributes to $N*D$ times LLM calling, where $D$ is the depth (or equivalently length) of the reasoning path, and $N$ is the width of the beam-search (how many paths are remained in the pool in each iteration). Specifically, we need to call $ND+D+C$ times LLM for each sample question, where $C$ is a constant (equals to $1$ if there is no error occurs when calling the API). [38] illustrates that the computational efficiency can be alleviated by replacing LLMs with small models such as BM25 and Sentence-BERT for the beam search decision since the small models are much faster than LLM calling. In this way, we can reduce the number of LLM calling from $ND+D+C$ to $D+C$. However, [38] shows that this optimization sacrifices the accuracy due to the weaker scoring model in decision making.
Table 6: Average number of LLM calls per question.
Dataset WebQSP CWQ CR-LT
Average LLM Calls 10.7 15.2 17.3
@@
Table 7: Average token usage of LLM calls per question.
Dataset WebQSP CWQ CR-LT
Average Token Usage 2,452 2,741 1,824
Cost efficiency. Another benefit of FiDeLiS is the cost efficiency. We control the token usage for each LLM calling through Path-RAG module. Specifically, Path-RAG only retrieves the top-n reasoning steps as the candidates for each beam search decision making, before making calls to the LLM. This approach ensures token expenditure without compromising the model’s performance, especially in some cases where hundreds of neighbors exist around a entity node. Table 7 shows the average token usage per question needed by FiDeLiS on different datasets. It can be seen that in three KGQA datasets, the numbers of token usages are ranging from 1,824 to 2,751, which could significantly reduce the cost for running the method.
Appendix E Variants of Path-RAG
Variant-1: To enhance the scoring function for reasoning in a knowledge graph and ensure each entity and relation are consistently linked, we can modify the scoring function to more explicitly account for the connectivity and interdependences between entities and relations throughout the path. This can be done by incorporating an accumulative sum that keeps track of previous scores and emphasizes path completeness. The revised formula introduces variables for the accumulated scores of entities and relations, which will help ensure all links in the path are considered:
$S(p)=S_{r}^{i}+S_{e}^{i}+\alpha\max_{\forall j\in N_{i}}\left(S_{r}^{j}+S_{e}^{j}\right)+\beta\sum_{k=0}^{i-1}\left(S_{r}^{k}+S_{e}^{k}\right)$ (8)
where: (1) $S_{r}^{i}$ and $S_{e}^{i}$ denote the relevance score of the current step relation and the corresponding neighboring entity, respectively. (2) $S_{r}^{j}$ and $S_{e}^{j}$ denote the relevance score of the 1-hop relation and the corresponding neighboring entity from node $N_{i}$. (3) The new term, $\sum_{k=0}^{i-1}\left(S_{r}^{k}+S_{e}^{k}\right)$, accumulates the scores from all previous steps from the start up to the step before the current, enhancing the influence of the entire path on the decision of the next step. (4) $\alpha$ and $\beta$ are weight factors that control the influence of the maximum score of neighboring nodes and the accumulated score to adjust the balance in prioritizing extended paths versus immediate gains. This adjustment helps to: firstly, reward paths that maintain strong links throughout, discouraging breaks or weakly connected steps in the reasoning process; secondly, enhance the depth of consideration in the decision-making process, looking both at immediate and long-term implications of each step.
Variant-2: We also consider to incorporate the randomness into the scoring function to balance the quality-diversity trade-off in searching for better reasoning paths as follows:
$S(p)=\lambda_{1}\left(S_{r}^{i}+S_{e}^{i}+\alpha\sum_{j\in N_{i}}(S_{r}^{j}+S_{e}^{j})\right)+\lambda_{2}\sum_{k=0}^{i-1}\left(D(p,P_{k})\right)$ (9)
where $D(p,P_{\text{sel}})$ refers to the distance between the selected path with the grounded reasoning paths. while $\lambda_{1}$ and $\lambda_{2}$ are weighting factors that balance the importance of internal quality metrics against diversity metrics.
Appendix F Additional Related Works: Reasoning with LLM Prompting
End-to-End methods [4, 15] aims to leverage LLMs to generate final answers directly, often done by providing a task description and/or a few examples for in-context learning. While it offers convenience, however, it suffers from the un-interpretability of the generation, and the lack of explicit steps and reliance solely on the LLMs’ training data may result in un-robustness and exhibit sensitivity to slight input variations.
Chain-of-Thought methods [48, 23, 8] emphasize breaking down complex reasoning into a series of intermediate steps to support complex reasoning. However, CoT suffers from unreliability and uncontrollability since the generated reasoning steps may not always align with the intended logical progression or may produce incorrect or inconsistent answers when multiple questions are posed consecutively.
Plan-and-solve methods [45] prompts LLMs to generate a plan and conduct reasoning based on it. DecomP [12] prompts LLMs to decompose the reasoning task into a series of sub-tasks and solve them step by step.
Semantic Parsing/Code Generation methods [5, 53, 7] leverages LLMs to convert natural language queries into executable code or structured representations. This approach enables more precise control over the output and facilitates better interpretability. However, it has its limitations, such as the limited coverage of programming language grammar and semantics. This restricts the model’s ability to handle complex programming tasks and may require additional techniques to handle out-of-domain queries effectively. [5] propose to leverage LLMs to bridge the out-of-domain queries, however, the form of knowledge injection still lacks a thorough exploration and the trigger functions are still naive for complex queries. However, the problem of hallucinations and lack of knowledge affect the faithfulness of LLMs’ reasoning. ReACT [51] treats LLMs as agents, which interact with the environment to get the latest knowledge for reasoning. To explore faithful reasoning, FAME [16] introduces the Monte-Carlo planning to generate faithful reasoning steps. RR [13] and KD-CoT [44] further retrieve relevant knowledge from KGs to produce faithful reasoning plans for LLMs.
Appendix G Experiment Details
G.1 Baselines
RoG [29]: It aims to embed structure knowledge from KGs into the underling neural networks during the pretraining and fine-tuning process. However, the controllablity of the reasoning paths generation is hard to achieve since the decoding process can be shortsighted, i.e., it is difficult to tell whether the current operation/decision would lead to a malicious operation several steps later. Also, the reasoning steps generated by LLMs are observed to be prone to errors, which could be non-exist in the knowledge graph, and can subsequently lead to incorrect reasoning during inference. Additionally, KG embedded in LLM sacrifices its own nature of explainability in knowledge reasoning and efficiency in knowledge updating [17]. Compared to RoG, FiDeLiS propose to employ beam search tailored for intermediate reasoning steps and guide the beam searching process by controlling the error of each reasoning step to prevent accumulation thought the chaining.
ToG [38]: It aims step-wise grounding to elicit more accurate results. However, it directly employs LLMs to output scores for candidate selections from the KG, which require an LLM/LM to iteratively score candidate plans, which are resource-consuming in terms of both time and computing. Also, the performance of the proposed ToG [38] seems rely heavily on the use of very strong LLMs (including Codex [3], GPT-4 [31]) to do the final reasoning; and the true potential of enabling few-shot grounded reasoning path understanding has yet to be fully realized.
G.2 Datasets
@@
Motivation of CR-LT-KGQA The motivation for evaluating over CR-LT-KGQA is that the majority of existing KGQA datasets, including WebQSP and CWQ, predominantly feature popular entities. These entities are well-represented in the training corpora of LLMs, allowing to often generate correct answers based on their internal knowledge, potentially without external KGs. Moreover, since WebQSP and CWQ have been available for several years, there is a likelihood that their data might have been incorporated into the pre-training corpora of LLMs, further reducing the need for external KGs during question-answering. Against this backdrop, we specifically chose CR-LT-KGQA for our evaluation as it includes queries related to obscure or long-tail entities, the distribution of the frequency and entity popularity of CR-LT is shown as Figure 5. Such scenarios are where KGs play a critical role because they provide a reliable source of verifiable information, crucial when LLMs encounter entities not well covered in their training data. By testing our methods on CR-LT-KGQA, we aim to explore how effective LLMs can operate when combined with KGs, especially in contexts involving less common knowledge domains where LLM performance typically declines. This evaluation helps us understand the extent to which KGs remain necessary for supporting LLMs in a diverse range of query scenarios.
Table 10: Comparison of FiDeLiS using different backbone LLMs.
Backend Models Methods WebQSP CWQ CR-LT
Hits@1 (%) F1 (%) Hits@1 (%) F1 (%) Acc (%)
Llama-2-13B FiDeLiS 72.34 69.78 58.41 54.78 60.87
Mistral-7B FiDeLiS 74.11 70.23 60.71 56.87 63.12
Example of Deductive Verification Question: “Who is the ex-wife of Justin Bieber’s father?" After one round of beam searching, the current reasoning path is “Justin_bieber $\to$ people.person.father $\to$ Jeremy_bieber", and the next step candidates are “1. people.married_to.person $\to$ Erin Wagner"; 2. people.person.place_of_birth $\to$ US, $\ldots$". The deductive reasoning can be formulated as: Premises: Justin_bieber $\to$ people.person.father —> Jeremy_bieber (from the current reasoning path) Jeremy_bieber $\to$ people.married_to.person $\to$ Erin Wagner (from the next step candidates) Conclusion: Erin Wagner is the ex-wife of Justin Bieber’s father. (Using LLM zero-shot to reformat the question to a cloze filling, and use the last entity from the next step candidate"Erin Wagner” to fill the cloze) The prompt will be whether the conclusion can be deduced from the given premises, if yes, return “yes”, it no, return “no".
Appendix H Algorithm Description

### v2

@@ 5 Conclusion
Appendix A Definitions
A.1 Knowledge Graph-based Question Answering (KGQA)
@@
Knowledge Graph Question Answering (KGQA) is a typical reasoning task based on KGs. Given a natural language question $q$ and a $\mathrm{KG}$ $\mathcal{G}$, the task aims to design a function $f$ to predict answers $a\in\mathcal{A}_{q}$ based on knowledge from $\mathcal{G}$, i.e., $a=f(q,\mathcal{G})$. Following previous works Luo et al. (2024; 2023), we assume the entities $e_{q}\in\mathcal{T}_{q}$ mentioned in $q$ and answers $a\in\mathcal{A}_{q}$ are labeled and linked to the corresponding entities in $\mathcal{G}$, i.e., $\mathcal{T}_{q},\mathcal{A}_{q}\subseteq\mathcal{E}$.
Appendix B Potential Impacts and Limitations
@@
The proposed method holds the potential to significantly enhance the performance of large language models (LLMs) by tackling the issue of hallucinations, thereby fostering deep, responsible reasoning. By integrating KGs with LLMs, the approach not only facilitates more accurate knowledge retrieval but also leverages deductive reasoning capabilities to steer the reasoning process and circumvent logical fallacies. The method is characterized by its stepwise, generalizable approach, and the use of deductive verification as a stopping criterion, which together may reduce superfluous computation and curb misleading reasoning chains. Moreover, due to its training-free nature and lower computational demands, this method could seamlessly serve as a plug-in for other existing frameworks.
@@
However, it’s important to recognize certain limitations. Since the method is still in its early stages of development, it might encounter unanticipated challenges or drawbacks in real-world applications. Its dependency on external KGs means that the quality and comprehensiveness of these resources can affect its overall effectiveness. Computational constraints might also arise when dealing with very large or intricate graphs. Furthermore, although the method has demonstrated promise in benchmark tests, its performance across more diverse or specialized tasks remains untested.
Appendix C Bottleneck of Beam Search Efficiency
@@
The bottleneck of computation is the beam search process, which contributes to $N*D$ times LLM calling, where $D$ is the depth (or equivalently length) of the reasoning path, and $N$ is the width of the beam-search (how many paths are remained in the pool in each iteration). Specifically, we need to call $ND+D+C$ times LLM for each sample question, where $C$ is a constant (equals to $1$ if there is no error occurs when calling the API). Sun et al. (2023) illustrates that the computational efficiency can be alleviated by replacing LLMs with small models such as BM25 and Sentence-BERT for the beam search decision since the small models are much faster than LLM calling. In this way, we can reduce the number of LLM calling from $ND+D+C$ to $D+C$. However, Sun et al. (2023) shows that this optimization sacrifices the accuracy due to the weaker scoring model in decision making.
@@
Appendix D Experiment Details
D.1 Baselines
Baselines. (1) RoG (Luo et al., 2024): embed structure knowledge graph from KGs into the underlying neural networks during the pretraining and fine-tuning process to generate the reasoning path and explanation. (2) ToG (Sun et al., 2023): ask LLM to iteratively explore multiple possible reasoning paths on KGs until the LLM determines that the question can be answered based on the current reasoning paths. Our method FiDeLiS follows the similar paradigm. (3) NSM (He et al., 2021) utilizes the sequential model to mimic the multi-hop reasoning process. (4) KD-CoT (Wang et al., 2023a) retrieves relevant knowledge from KGs to generate faithful reasoning paths for LLMs. (5) DeCAF (Yu et al., 2022) combines semantic parsing and LLMs reasoning to jointly generate answers, which also reach salient performance on KGQA tasks.
Implementation Details. We set the default beam width as $4$ and depth as $4$ without specific annotation. We set the $\alpha$ in Eq 3 as $0.3$ to ensure reproducibility. For LLMs usage, we set all the inference using temperature $T=0.3$ and $p=1.0$. The hyperparameter tuning experiments for beam search and $\alpha$ can be found in Figure 2 and Table 9.
Backboned LLMs. We assess our approach on closed- and open-source LLMs. For closed-source LLMs, we choose GPT-4-turbo, GPT-3.5-turbo [^1] We use the recent gpt-4-turbo model that is released in 2024-04-09 from https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4 and gpt-3.5-turbo model notated as “gpt-3.5-turbo-0125” from https://platform.openai.com/docs/models/gpt-3-5-turbo. to report and compare the results on all datasets. We use Llama-2-13B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023a) as our open-source LLMs to conduct cost–performance analysis on different datasets. The experiment results of open-source models can be found in Table 10.
@@ D.1 Baselines
D.2 Datasets
Datasets & Metrics. We adopt three benchmark KGQA datasets: WebQuestionSP (WebQSP) (Yih et al., 2016), Complex WebQuestions (CWQ) (Talmor & Berant, 2018) and CR-LT-KGQA (Guo et al., 2024) in this work. We follow previous work (Luo et al., 2024) to use the same training and testing splits for fair comparison over WebQSP and CWQ. The questins from both WebQSP and CWQ can be reasoned using Freebase KGs [^2] https://github.com/microsoft/FastRDFStore. To address the bias in WebQSP and CWQ, which predominantly feature popular entities and there is a likelihood that their data might have been incorporated into the pre-training corpora of LLMs, we further test our method on CR-LT-KGQA (further discussed in Appendix Section 8). We use the complete dataset from CR-LT-KGQA in our experiments, as it comprises only 200 samples. Each of the question can be reasoned based on the Wikidata [^3] https://www.wikidata.org/wiki/Wikidata:Main_Page. The statistics of the datasets are given in Table 7 and Table 8. To streamline the KGs, we utilize a subgraph of Freebase by extracting all triples that fall within the maximum reasoning hops from the question entities in WebQSP and CWQ followsing RoG (Luo et al., 2024). Similarly, we construct the corresponding sub-graphs of Wikidata for CR-LT-KGQA as well. We assess the performance of the methods by analyzing the F1 and Hits@1 metrics for CWQ and WebQSP, and by evaluating the accuracy for CR-LT-KGQA.
@@ D.2 Datasets
Motivation of CR-LT-KGQA. The motivation for evaluating over CR-LT-KGQA is that the majority of existing KGQA datasets, including WebQSP and CWQ, predominantly feature popular entities. These entities are well-represented in the training corpora of LLMs, allowing to often generate correct answers based on their internal knowledge, potentially without external KGs. Moreover, since WebQSP and CWQ have been available for several years, there is a likelihood that their data might have been incorporated into the pre-training corpora of LLMs, further reducing the need for external KGs during question-answering.
@@
Against this backdrop, we specifically chose CR-LT-KGQA for our evaluation as it includes queries related to obscure or long-tail entities, the distribution of the frequency and entity popularity of CR-LT is shown as Figure 4. Such scenarios are where KGs play a critical role because they provide a reliable source of verifiable information, crucial when LLMs encounter entities not well covered in their training data. By testing our methods on CR-LT-KGQA, we aim to explore how effective LLMs can operate when combined with KGs, especially in contexts involving less common knowledge domains where LLM performance typically declines. This evaluation helps us understand the extent to which KGs remain necessary for supporting LLMs in a diverse range of query scenarios.
D.3 Parameter Tuning for $\alpha$ for Scoring Function
For hyperparameter tuning regarding $\alpha$ for Eq 3, we added an extra comparison in Table 9 which shows that $\alpha$ is actually not a very impact parameter for the whole system, however for the reproducibility of our method, we set the $\alpha$ as 0.3.
Table 9: Parameter tuning for $\alpha$ for scoring function over WebQSP
$\alpha$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0
Hits@1 83.76 84.15 84.39 83.98 83.55 84.12 83.79 82.25 83.60 83.73
Appendix E Algorithm Description
@@
Appendix F Prompt List
F.1 Plan-and-solve
In-Context Few-shot
Q: {Query}
A:
F.2 Deductive-verification
In-Context Few-shot
A:
F.3 Adequacy-verification
In-Context Few-shot
A:
F.4 Beam Search
In-Context Few-shot
A:
F.5 Reasoning
In-Context Few-shot
Question: {Query}
Reasoning path: {reasoning_path}
Only return the answer to the question.
A:
F.6 Demonstration of Deductive Verification
Deductive Verification Example Question: Who is the ex-wife of Justin Bieber’s father? After one round of beam searching, the current reasoning path is: Justin_bieber $\to$ people.person.father $\to$ Jeremy_bieber. The next step candidates are: 1. people.married_to.person $\to$ Erin Wagner 2. people.person.place_of_birth $\to$ US, $\ldots$ The deductive reasoning can be formulated as follows: Premises: - Justin_bieber $\to$ people.person.father $\to$ Jeremy_bieber (from the current reasoning path) - Jeremy_bieber $\to$ people.married_to.person $\to$ Erin Wagner (from the next step candidates) Conclusion: Erin Wagner is the ex-wife of Justin Bieber’s father. (Using a large language model (LLM) zero-shot approach to reformat the question into a cloze filling task, we use the last entity from the next step candidates, "Erin Wagner", to fill the cloze.) The prompt will ask whether the conclusion can be deduced from the given premises. If the answer is "yes", return “yes”, otherwise return “no.”
Table 10: Comparison of FiDeLiS using different backbone LLMs.
Backend Models Methods WebQSP CWQ CR-LT
Hits@1 (%) F1 (%) Hits@1 (%) F1 (%) Acc (%)
Llama-2-13B FiDeLiS 72.34 69.78 58.41 54.78 60.87
Mistral-7B FiDeLiS 74.11 70.23 60.71 56.87 63.12
Appendix G Related Works: Reasoning with LLM Prompting
End-to-End methods (Chen, 2022; Hoffmann et al., 2022) aims to leverage LLMs to generate final answers directly, often done by providing a task description and/or a few examples for in-context learning. While it offers convenience, however, it suffers from the un-interpretability of the generation, and the lack of explicit steps and reliance solely on the LLMs’ training data may result in un-robustness and exhibit sensitivity to slight input variations.
Chain-of-Thought methods (Wei et al., 2022a; Kojima et al., 2022; Gong et al., 2020) emphasize breaking down complex reasoning into a series of intermediate steps to support complex reasoning. However, CoT suffers from unreliability and uncontrollability since the generated reasoning steps may not always align with the intended logical progression or may produce incorrect or inconsistent answers when multiple questions are posed consecutively.
Plan-and-solve methods (Wang et al., 2023b) prompts LLMs to generate a plan and conduct reasoning based on it. DecomP He et al. (2021) prompts LLMs to decompose the reasoning task into a series of sub-tasks and solve them step by step.
Semantic Parsing/Code Generation methods (Cheng et al., 2023; Ye et al., 2023; Gemmell & Dalton, 2023) leverages LLMs to convert natural language queries into executable code or structured representations. This approach enables more precise control over the output and facilitates better interpretability. However, it has its limitations, such as the limited coverage of programming language grammar and semantics. This restricts the model’s ability to handle complex programming tasks and may require additional techniques to handle out-of-domain queries effectively. Cheng et al. (2023) propose to leverage LLMs to bridge the out-of-domain queries, however, the form of knowledge injection still lacks a thorough exploration and the trigger functions are still naive for complex queries. However, the problem of hallucinations and lack of knowledge affect the faithfulness of LLMs’ reasoning. ReACT Yao et al. (2022) treats LLMs as agents, which interact with the environment to get the latest knowledge for reasoning. To explore faithful reasoning, FAME Hong et al. (2023) introduces the Monte-Carlo planning to generate faithful reasoning steps. RR He et al. (2022) and KD-CoT Wang et al. (2023a) further retrieve relevant knowledge from KGs to produce faithful reasoning plans for LLMs.